OpenStudy (anonymous):

If the instructions for a problem ask you to use the smallest possible domain to completely graph two periods of y = 5 + 3 cos 2(x - pi/3), what should be used for Xmin and Xmax? Explain your answer.

OpenStudy (sirm3d):

a simple solution to find the \(x_{\text{min}}\) and \( x_{\text{max}}\) for one period, solve the two equations separately: \[x-\frac{\pi}{3}=0\\x-\frac{\pi}{3}=k\pi\] for one complete period, \(k=1\). one of the solutions is \(x_{\text{min}}\), while the other solution is \(x_{\text{max}}\)

OpenStudy (anonymous):

x = pi/3, x = pi - (pi/3) ?

OpenStudy (anonymous):

this one, @Shane_B . thanks for your time.

OpenStudy (shane_b):

I believe you have it right :)

OpenStudy (anonymous):

lol, how? i don't get it @_@

OpenStudy (shane_b):

I wrote out a long and drawn out answer but I didn't like it...give me another min or two :)

OpenStudy (shane_b):

I hope this makes sense to you: The range of cos is from -1 to 1. The values at -1 and 1 will be the same except for the sign. Therefore, you just need to find the values of x that would equal cos(0) and the values of x that would equal cos(pi). That will give you half a wave but the second half would only be the inverse of it.

OpenStudy (anonymous):

"That will give you half a wave but the second half would only be the inverse of it" is where you lost me. sorry. hehe but thanks. I appreciate the help, I'll try to make sense out of it. ><

OpenStudy (shane_b):

A full cos/sin wave is 0 to 2pi...a half wave is 0 to pi. :) Maybe this will make it clearer:

OpenStudy (anonymous):

oh. yeah. ^^'

OpenStudy (shane_b):

Note that the second part of the wave (the part that falls below the line) is the inverse of the first part. That's how sin and cos waves always look :)

OpenStudy (anonymous):

So, to get the first part of the answer, I need to find the values of x that would equal cos(0) and the values of x that would equal cos(pi)?

OpenStudy (anonymous):

pi/2, 3pi/2.... I don't know what's cos(pi).

OpenStudy (shane_b):

\[cos(\pi)\]

OpenStudy (shane_b):

Remember @sirm3d's response? Basically you just solve these two equations: \[x-\frac{\pi}{3}=0\\x-\frac{\pi}{3}=\pi\]So you end up with the following for min/max:\[x=\frac{\pi}{3}\]\[x=\frac{2}{3}\pi\] You already solved these btw :)

OpenStudy (anonymous):

^^' I see it's more simplified in your mind. I think my biggest problem is how to explain this answer to my teacher...

OpenStudy (shane_b):

Yea, I'm not doing very well at explaining this either. Maybe someone else can put it into words better :/

OpenStudy (shane_b):

I think I'd just explain with a picture :)